# Federated Unsupervised Domain Generalization using Global and Local Alignment of Gradients

### v1

@@ A.2 Proof of Claim 1
$g=\left\{\begin{array}[]{ll}(\sigma(Wx_{1}-Wx_{2})-1)(x_{1}-x_{2})&\text{if }y=1\text{ (Positive pairs)}\\ \sigma(Wx_{1}-Wx_{2})(x_{1}-x_{2})&\text{if }y=0\text{ (Negative pairs)}\end{array}\right.$ (10)
@@
$\frac{\partial{g^{+}}}{\partial x_{1}}\bigg{|}_{x_{1}=\mu_{1}}$ $=\sigma(W\mu_{1}-WA\mu-WB)(1-\sigma(W\mu_{1}-WA\mu_{1}-WB))(\mu_{1}-A\mu-B)$
$+\sigma(W\mu_{1}-WA\mu_{1}-WB)(1-A).$ (12)
@@
$\frac{\partial{g^{+}}}{\partial x_{1}}\bigg{|}_{x_{1}=\mu_{1}}=\sigma(-WB)(1-\sigma(-WB))(-B)+\sigma(-WB)(1-A).$ (13)
@@
$\frac{\partial{g^{-}}}{\partial x_{1}}\bigg{|}_{x_{1}=\mu_{1}}$ $=\sigma(W\mu_{1}-Wx_{2})(1-\sigma(W\mu_{1}-Wx_{2}))(\mu_{1}-Ax_{2})$
$+\sigma(W\mu_{1}-Wx_{2})(1-x_{2}).$ (15)
@@
$\frac{\partial{g^{-}}}{\partial x_{1}}\bigg{|}_{x_{1}=\mu_{1}}=\sigma(-Wx_{2})(1-\sigma(-Wx_{2}))(-Ax_{2})+\sigma(-Wx_{2})(1-x_{2}).$ (16)
@@
$\frac{\partial g^{-}}{\partial x_{1}}\bigg{|}_{x_{1}=\mu_{1}}$ $\simeq\left(\frac{1}{2}+\sum_{n=0}^{\infty}\frac{(-1)^{n+1}(2^{2n+1}-1)B_{2n+2}}{(2n+2)!}(Wx_{2})^{2n+1}\right)$
×$\quad\times\left(\frac{1}{2}-\sum_{n=0}^{\infty}\frac{(-1)^{n+1}(2^{2n+1}-1)B_{2n+2}}{(2n+2)!}(Wx_{2})^{2n+1}\right)(-Ax_{2})$
$\quad+\left(\frac{1}{2}+\sum_{n=0}^{\infty}\frac{(-1)^{n+1}(2^{2n+1}-1)B_{2n+2}}{(2n+2)!}(Wx_{2})^{2n+1}\right)(1-x_{2}).$ (18)
Since the coeffiencts are only consist of $n$, $B_{n}$ and $W$, we can simplify as:
@@
By applying this estimate to Eq. ( 20), we derive:
@@
Consequently, we can conclude that for all negative pairs, the gradient derivative $\frac{\partial{g^{-}}}{\partial x_{i}}\bigg{|}_{x_{i}=\mu_{i}}$ only depends on the network weights. Hence, $\text{sign}(\frac{\partial{g}_{i}}{\partial x^{f}_{i}}\big{|}_{x^{f}_{i}=\mu^{f}_{i}})=\text{sign}(\frac{\partial{g}_{j}}{\partial x^{f}_{j}}\big{|}_{x^{f}_{j}=\mu^{f}_{j}}).$
A.3 Proof of Corollary 1
Proof.
The variance of $\textbf{g}_{i}$ and $\textbf{g}_{j}$ is:
Lemma A1.
[…]
A.4 Proof of Lemma A1
Proof.
The variance of g, by definition, is:
The squared term can be expanded as:
$\text{Var}(\textbf{g})=\mathbb{E}\left[\sum_{f=1}^{F}\sum_{e=1}^{F}(x^{f}-\mu^{f})(x^{e}-\mu^{e})\left(\frac{\partial g}{\partial x^{f}}\bigg{|}_{x^{f}=\mu^{f}}\right)\right.\left.\left(\frac{\partial g}{\partial x^{e}}\bigg{|}_{x^{e}=\mu^{e}}\right)\right].$ (30)
$\text{Var}(\textbf{g})=\mathbb{E}\left[\sum_{f=1}^{F}(x^{f}-\mu^{f})^{2}\left(\frac{\partial g}{\partial x^{f}}\bigg{|}_{x^{f}=\mu^{f}}\right)^{2}\right].$ (31)
Since $\mathbb{E}{(x^{f}-\mu^{f})^{2}}=(\sigma^{f})^{2}$, we derive:
which completes the proof for the variance of g. ∎
A.5 Federated supervised domain generalization
@@ A.6 Proof of Claim A1
Appendix B Algorithm
Algorithm 1 lays out the detailed procedure of our proposed method.
Appendix C Self-supervised learning frameworks
In this study, we employ four SSL frameworks at the client level. These techniques include SimCLR [36], MoCo [50], BYOL [51], and SimSiam [55]. Each of these models consists of two encoders. BYOL and MoCo utilize exponential moving averages to update one of the encoders, namely the target network, using the other encoder, referred to as the online network. In contrast, SimCLR and SimSiam share weights between their two encoders. Additionally, SimCLR and MoCo, being contrastive-based SSL models, leverage negative samples in the learning process. Conversely, BYOL and SimSiam, which do not rely on negative samples, include a predictor on top of the online encoder to enhance learning from unlabeled samples.
Appendix D Implementation details
@@
Algorithm 1 FedGaLA
1:Input: data $\mathcal{D}_{i}$, initialization $\Theta^{0}$
2:Output: $\Theta^{T}$
[…]
4:     Server:
6:           Initialize global update: $\textbf{g}^{(t+1)}\leftarrow\text{FedAvg}(\textbf{g}_{i}^{(t)}$
[…]
9:               Normalize weights: $w_{i}\leftarrow\frac{w_{i}}{\sum w_{i}}$
10:               Aggregate updates: $\textbf{g}^{(t+1)}\leftarrow\frac{1}{K}\sum_{i=1}^{K}w_{i}\textbf{g}_{i}^{(t)}$
11:     end for
13:           Communicate: $\Theta_{i}\leftarrow\Theta^{(t+1)}$
14:     Client:
[…]
20:                  Update weights: $\theta_{i}^{(l,t)}\leftarrow\theta_{i}^{(l,t-1)}-\eta\cdot\nabla l_{i}^{(l,t)}(x_{i,j},\theta_{i}^{(l,t)})$
21:              end if
22:         end for
23:     end for
24:           Communicate: $S\leftarrow\Theta_{i}$
25:end for
D.1 DomainNet classes
Following [45], for the DomainNet dataset, we select the following classes: zigzag, tiger, tornado, flower, giraffe, toaster, hexagon, watermelon, grass, hamburger, blueberry, violin, fish, sun, broccoli, Eiffel tower, horse, train, bird, and bee [45], which results in a total of 38556 samples. In all experiments using this dataset, three domains are used for training (‘Painting’, ‘Real’, and ‘Sketch’) and the model is tested using the other three domains (‘Clipart’, ‘Infographics’, and ‘Quickdraw’).
D.2 Image augmentations
We follow [36] for augmentations. For all datasets, a random patch of the image is selected and resized to $32\times 32$. Subsequently, we apply two random transformations, namely horizontal flip and color distortion.
D.3 Network architecture
Predictor. For BYOL, we use a two-layer multilayer perceptron as the predictor. It begins with a fully connected layer featuring 4096 neurons, followed by one-dimensional batch normalization and a ReLU activation function. It concludes with another fully connected layer comprising 2048 neurons.
Encoder. We use ResNet18 [53] as the encoder for all the experiments. We use the ResNet architecture presented in [27], which is slightly different from the original architecture: (i) The first convolution layer employs a $3\times 3$ kernel size, replacing the original $7\times 7$; (ii) An average pooling layer with a $4\times 4$ kernel size is used before the final linear layer, substituting the adaptive average pooling layer; and (iii) The last linear layer is replaced with a two-layer MLP, which shares the same network architecture as the predictor.
D.4 Training and evaluation
@@
Training. While training the clients, we use a batch size of 128. Each client is trained for 7 local epochs before being returned to the server for a communication round. By default, we train for 100 communication rounds. We use the Adam [56] optimizer with a learning rate of $3\times 10^{-3}$. The hyperparameters used for training FedSimCLR are the same as those used in FedGaLA. For other baselines (FedMoCo, FedSimSiam, FedBYOL, and FedEMA), we use the SGD optimizer with a momentum of 0.9 and a weight decay of $3\times 10^{-4}$. The choice of learning rate for FedSimSiam, FedBYOL, and FedEMA is $0.03$ while for FedMoCo we use a learning rate of $0.025$. The parameters have been tuned to maximize performance.
@@
D.5 Regularizers
Below we provide the details of the L2-norm and proximal term regularizers used in [9] and [16], respectively.
@@ D.5 Regularizers
where $\lambda$ is the the regularization coefficient and $\|\Theta_{i}\|_{2}^{2}$ is the square of the L2-norm of the parameters, defined as:
@@
Proximal term. We also utilize the proximal term from [16] to further penalize the deviation of local models from the global model. The formulation of the proximal term is
@@
Appendix E Additional results and analysis
FedGaLA with other SSL frameworks. Tabel 5 illustrates the performance of FedGaLA when employing various SSL models in local training and compares its performance with the baseline. These findings highlight FedGaLA’s capability to boost the performance of different SSL techniques when used for federated unsupervised domain generalization.
Models P A C S Ave. FedMoCo 46.2(1.1) 28.6(1.0) 46.7(0.6) 37.6(1.3) 39.8 FedGaLA(MoCo) 46.5(0.3) 28.9(0.4) 47.9(0.3) 39.6(2.4) 40.7 FedBYOL 52.1(1.1) 31.8(1.1) 45.4(2.2) 47.4(1.9) 44.2 FedGaLA(BYOL) 52.8(0.6) 31.7(0.5) 46.3(3.1) 47.6(1.2) 44.6 FedSimCLR 64.2(1.2) 41.9(1.5) 58.4(1.3) 70.1(1.2) 58.6 FedGaLA(SimCLR) 64.7(1.9) 44.2(1.2) 60.5(2.2) 70.5(1.3) 60.0
Batch size. Figure 4(a) illustrates the impact of batch size on the performance of the network. Our experiment demonstrates that a batch size of 128 produces the best results.
Global gradient alignment iterations. Figure 4(b) shows the impact of different gradient alignment iterations (line 7 in Algorithm 1) on performance. FedGaLA produces the best results when the iteration number is set to 3. When the number of iterations increases the model places too much emphasis on clients that are aligned with the global average and decreases the impact of unaligned clients harshly, which leads to a decline in the overall performance of the model. On the other hand, with small iteration values, global gradient alignment is not performed effectively, and thus performance declines.
Figure 4: Impact of batch size (left) and the number of iterations for global alignment (right) on performance.
FedGaLA vs. FedAVG. In this experiment, we further evaluate the effectiveness of FedGaLA. To achieve this, after completing the aggregation process of FedGaLA, the global model is further updated using a combination of the weighted aggregated model and the FedAvg of the client models. This is formulated as:
$\theta^{t+1}\leftarrow\gamma(\theta^{t}+\hat{\textbf{g}}^{t+1})+(1-\gamma)\text{FedAvg}(\theta_{i}^{t+1}),$ (41)
where $\gamma$ is a parameter that balances the influence of the current aggregated update $\textbf{g}^{t+1}$ and the global model parameters obtained through FedAvg. FedAvg can be defined as $\theta^{t+1}=\theta^{t}+\frac{1}{N}\sum_{i=1}^{N}\tilde{w}_{i}\hat{\textbf{g}}_{i}^{t}$, where $\tilde{w}_{i}$ are the client weights, adjusted uniformly.
Table 6 shows the impact of $\gamma$ (introduced in Eq. (41) on performance. We observe that FedGaLA, i.e., $\gamma=1$, achieves the best results for all domains, except for the ‘Photo’ domain, where $\gamma=0.75$ yields the highest accuracy. It should be noted that with $\gamma=0.00$ the global alignment step of FedGaLA is effectively not performed, yielding the same performance observed in the second row of Table 3.
Table 6: Impact of $\gamma$ on performance.
$\gamma$ P A C S Ave.
1.00 64.7(1.9) 44.2(1.2) 60.5(2.2) 70.5(1.3) 60.0
0.75 65.4(1.7) 41.4(1.6) 59.8(2.4) 69.4(0.7) 58.9
0.50 64.9(1.3) 42.4(1.2) 59.2(1.9) 69.8(0.4) 59.1
0.25 64.8(1.2) 42.8(1.1) 59.7(1.2) 69.9(0.4) 59.3
0.00 64.7(0.4) 42.6(1.2) 58.1(0.6) 69.9(1.1) 58.8
Appendix F Additional discussion
Cross-silo vs. cross-device federated learning. Two frameworks can generally be considered in federated learning: cross-silo and cross-device. In cross-silo, the number of clients $K$ is small, and each client is capable of retaining its local model and state across training rounds. In contrast, cross-device federated learning works under the assumption of having millions of stateless clients, where only a limited number of clients participate in training during each communication round. However, due to limitations in experimental settings, the majority of studies typically conduct experiments involving at most hundreds of clients [57, 27]. In this study, given the limited number of domains FedGaLA works within the cross-silo framework. Moreover, each client is equipped with the necessary resources to participate in the collaborative learning process. Additionally, we ensure that every client updates its local model with a sufficient number of epochs.
Broader impact. This paper aims to advance the field of federated learning under realistic conditions of domain shift among clients and the absence of labeled data. While there are various potential societal implications of our work, we believe that the most notable is the advancement of decentralized privacy-preserving representation learning. Our approach enhances model generalization to new domains by aligning gradients across clients and the server. This leads to more accurate and robust models in areas such as the healthcare industry, where privacy is critical.

### v2

@@ A.3 Proof of Claim 1
$g=\\ \left\{\begin{array}[]{ll}(\sigma(Wx_{1}-Wx_{2})-1)(x_{1}-x_{2})&\text{if }y=1\\ &\text{ (Positive pairs)}\\ \sigma(Wx_{1}-Wx_{2})(x_{1}-x_{2})&\text{if }y=0\\ &\text{ (Negative pairs)}\\ \end{array}\right.$ (20)
@@
$\frac{\partial{g^{+}}}{\partial x_{1}}\bigg{|}_{x_{1}=\mu_{1}}=\\ \sigma(W\mu_{1}-WA\mu-WB)(1-\sigma(W\mu_{1}-WA\mu_{1}-WB))\\ (\mu_{1}-A\mu-B)+\sigma(W\mu_{1}-WA\mu_{1}-WB)(1-A).$ (22)
@@
$\frac{\partial{g^{+}}}{\partial x_{1}}\bigg{|}_{x_{1}=\mu_{1}}=\\ \sigma(-WB)(1-\sigma(-WB))(-B)+\sigma(-WB)(1-A).$ (23)
@@
$\frac{\partial{g^{-}}}{\partial x_{1}}\bigg{|}_{x_{1}=\mu_{1}}=\sigma(W\mu_{1}-Wx_{2})(1-\sigma(W\mu_{1}-Wx_{2}))\\ (\mu_{1}-Ax_{2})+\sigma(W\mu_{1}-Wx_{2})(1-x_{2}).$ (25)
@@
$\frac{\partial{g^{-}}}{\partial x_{1}}\bigg{|}_{x_{1}=\mu_{1}}=\sigma(-Wx_{2})(1-\sigma(-Wx_{2}))\\ (-Ax_{2})+\sigma(-Wx_{2})(1-x_{2}).$ (26)
@@
$\frac{\partial g^{-}}{\partial x_{1}}\bigg{|}_{x_{1}=\mu_{1}}\simeq\\ \left(\frac{1}{2}+\sum_{n=0}^{\infty}\frac{(-1)^{n+1}(2^{2n+1}-1)B_{2n+2}}{(2n+2)!}(Wx_{2})^{2n+1}\right)\quad\times\\ \left(\frac{1}{2}-\sum_{n=0}^{\infty}\frac{(-1)^{n+1}(2^{2n+1}-1)B_{2n+2}}{(2n+2)!}(Wx_{2})^{2n+1}\right)(-Ax_{2})\\ \quad+\left(\frac{1}{2}+\sum_{n=0}^{\infty}\frac{(-1)^{n+1}(2^{2n+1}-1)B_{2n+2}}{(2n+2)!}(Wx_{2})^{2n+1}\right)(1-x_{2}).\$ (28)
Since the coefficients are only consist of $n$, $B_{n}$ and $W$, we can simplify as:
@@
By applying this estimate to Eq. ( 30), we derive:
@@
Consequently, we can conclude that for all negative pairs, the gradient derivative $\frac{\partial{g^{-}}}{\partial x_{i}}\bigg{|}_{x_{i}=\mu_{i}}$ only depends on the network weights. Hence, due to the fact that both matrices are symmetric, $\text{sign}(\frac{\partial{g}_{i}}{\partial x^{f}_{i}}\big{|}_{x^{f}_{i}=\mu^{f}_{i}})=\text{sign}(\frac{\partial{g}_{j}}{\partial x^{f}_{j}}\big{|}_{x^{f}_{j}=\mu^{f}_{j}})$.
A.4 Federated supervised domain generalization
@@ A.5 Proof of Claim A1
A.6 Proof of Corollary 1
Proof.
The variance of $\textbf{g}_{i}$ and $\textbf{g}_{j}$ is:
Lemma A1.
$\text{Var}(\textbf{g}_{i}-\textbf{g}_{j})=\\ \sum_{f=1}^{F}(\sigma^{f}_{i})^{2}(\frac{\partial{g}_{i}}{\partial x^{f}_{i}}\bigg{|}_{x^{f}_{i}=\mu^{f}_{i}})^{2}+(\sigma^{f}_{j})^{2}(\frac{\partial{g}_{j}}{\partial x^{f}_{j}}\bigg{|}_{x^{f}_{j}=\mu^{f}_{j}})^{2}\\ -2\sigma_{x_{i}^{f},x_{j}^{f}}(\frac{\partial{g}_{i}}{\partial x^{f}_{i}}\bigg{|}_{x^{f}_{i}=\mu^{f}_{i}})(\frac{\partial{g}_{j}}{\partial x^{f}_{j}}\bigg{|}_{x^{f}_{j}=\mu^{f}_{j}}).$ (41)
A.7 Proof of Lemma A1
Proof.
The variance of g, by definition, is:
The squared term can be expanded as:
$\text{Var}(\textbf{g})=\\ \mathbb{E}[\sum_{f=1}^{F}\sum_{e=1}^{F}(x^{f}-\mu^{f})(x^{e}-\mu^{e})\left(\frac{\partial g}{\partial x^{f}}\bigg{|}_{x^{f}=\mu^{f}}\right)\\ \left(\frac{\partial g}{\partial x^{e}}\bigg{|}_{x^{e}=\mu^{e}}\right)].$ (45)
$\text{Var}(\textbf{g})=\\ \mathbb{E}\left[\sum_{f=1}^{F}(x^{f}-\mu^{f})^{2}\left(\frac{\partial g}{\partial x^{f}}\bigg{|}_{x^{f}=\mu^{f}}\right)^{2}\right].$ (46)
Since $\mathbb{E}{(x^{f}-\mu^{f})^{2}}=(\sigma^{f})^{2}$, we derive:
which completes the proof for the variance of g. ∎
A.8 Proof of Proposition 1
Assumption A2.
Let $\textbf{\emph{g}}_{i}$ and $\textbf{\emph{g}}_{j}$ be two sets of gradients. $\textbf{\emph{g}}_{est}$ is defined as the average of both sets. We assume $cos(\textbf{\emph{g}}_{j,k},\textbf{\emph{g}}_{est})>0$ for all $k\neq K$, $cos(\textbf{\emph{g}}_{i,k},\textbf{\emph{g}}_{est})>0$ for all $k$, and $cos(\textbf{\emph{g}}_{j,K},\textbf{\emph{g}}_{est})<0$. Without the loss of generality, we assume the last vector in the set is the unaligned vector.
Proof.
We first define $\hat{\textbf{g}_{j}}$ as:
$\textbf{g}_{j}=\hat{\textbf{g}_{j}}\cup\{\textbf{g}_{j,K}\}.$ (48)
Hence, the mean of $\textbf{g}_{j}$ can be derived as:
$\mu_{\textbf{g}_{j}}=\\ \frac{1}{K}((K-1)\mu_{\hat{\textbf{g}_{j}}}+\textbf{g}_{j,K}),\quad\mu_{\hat{\textbf{g}_{j}}}=\frac{1}{K-1}\sum_{k=1}^{K-1}\textbf{g}_{j,K}.$ (49)
Now we shift to the covariance of the two sets:
$\text{Cov}(\textbf{g}_{i},\textbf{g}_{j})=\frac{1}{K-1}(\textbf{g}_{i}-\mu_{\textbf{g}_{i}})(\textbf{g}_{j}-\mu_{\textbf{g}_{i}})^{T}.$ (50)
Substituting $\text{Cov}(\textbf{g}_{i},\hat{\textbf{g}_{j}})$ in Eq. 50, we derive:
$\text{Cov}(\textbf{g}_{i},\textbf{g}_{j})=\\ \frac{K-2}{K-1}\text{Cov}(\textbf{g}_{i},\hat{\textbf{g}_{j}})+\frac{1}{K-1}(\textbf{g}_{i}-\mu_{\textbf{g}_{i}})(\textbf{g}_{j,k}-\mu_{\textbf{g}_{i}})^{T}.$ (51)
By substituting $\mu_{\textbf{g}_{j}}$ in this equation we have:
$\text{Cov}(\textbf{g}_{i},\textbf{g}_{j})=\\ \frac{K-2}{K-1}\text{Cov}(\textbf{g}_{i},\hat{\textbf{g}_{j}})+\frac{1}{K-1}(\textbf{g}_{i}-\mu_{\textbf{g}_{i}})\\ (\textbf{g}_{j,k}-\frac{1}{K}((K-1)\mu_{\hat{\textbf{g}_{j}}}+\textbf{g}_{j,K}))^{T}.$ (52)
Simplifying this equation we derive:
$\text{Cov}(\textbf{g}_{i},\textbf{g}_{j})=\\ \frac{K-2}{K-1}\text{Cov}(\textbf{g}_{i},\hat{\textbf{g}_{j}})+\frac{1}{K}(\textbf{g}_{i}-\mu_{\textbf{g}_{i}})(\textbf{g}_{j,K}-\mu_{\hat{\textbf{g}_{j}}})^{T}.$ (53)
The term $(\textbf{g}_{i}-\mu_{\textbf{g}_{i}})(\textbf{g}_{j,K}-\mu_{\hat{\textbf{g}_{j}}})^{T}$ is a matrix with mostly negative components because of the following reasons. $(\textbf{g}_{i}-\mu_{\textbf{g}_{i}})$ represents the deviations of $\textbf{g}_{i}$ vectors from their mean which points around $\textbf{g}_{est}$ as $cos(\textbf{g}_{i,k},\textbf{g}_{est})>0$ for all $k$. On the other hand, since $cos(\textbf{g}_{j,K},\textbf{g}_{est})<0$, $(\textbf{g}_{j,K}-\mu_{\hat{\textbf{g}_{j}}})^{T}$ representing the deviation of $\textbf{g}_{j,K}$ from the mean vector is in the opposite direction of $\textbf{g}_{est}$. Hence, the dot product of two sets of vectors pointing in two opposite directions results in a negative covariance matrix which completes the proof. ∎
Appendix B Implementation details
@@
B.1 Image augmentations
We follow (Chen et al. 2020) for augmentations. For all datasets, a random patch of the image is selected and resized to $32\times 32$. Subsequently, we apply two random transformations, namely horizontal flip and color distortion.
B.2 Network architecture
We use ResNet18 (He et al. 2016) as the encoder for all the experiments. We use the ResNet architecture presented in (Zhuang, Wen, and Zhang 2022), which is slightly different from the original architecture: (i) The first convolution layer employs a $3\times 3$ kernel size, replacing the original $7\times 7$; (ii) An average pooling layer with a $4\times 4$ kernel size is used before the final linear layer, substituting the adaptive average pooling layer; and (iii) The last linear layer is replaced with a two-layer MLP, which shares the same network architecture as the predictor.
B.3 Training and evaluation
@@
Training. While training the clients, we use a batch size of 128. Each client is trained for 7 local epochs before being returned to the server for a communication round. By default, we train for 100 communication rounds. We use the Adam (Kingma and Ba 2014) optimizer with a learning rate of $3\times 10^{-3}$. The hyperparameters used for training FedSimCLR are the same as those used in FedGaLA. For other baselines (FedMoCo, FedSimSiam, FedBYOL, and FedEMA), we use the SGD optimizer with a momentum of 0.9 and a weight decay of $3\times 10^{-4}$. The choice of learning rate for FedSimSiam, FedBYOL, and FedEMA is $0.03$ while for FedMoCo we use a learning rate of $0.025$. The parameters have been tuned to maximize performance.
@@
B.4 Regularizers
Below we provide the details of the L2-norm and proximal term regularizers used in (Nguyen, Torr, and Lim 2022) and (Li et al. 2020), respectively.
@@ B.4 Regularizers
where $\lambda$ is the regularization coefficient and $\|\Theta_{i}\|_{2}^{2}$ is the square of the L2-norm of the parameters, defined as:
@@
Proximal term. We also utilize the proximal term from (Li et al. 2020) to further penalize the deviation of local models from the global model. The formulation of the proximal term is: